# Towards Efficient Real-Time Video Motion Transfer via Generative Time Series Modeling

### v1

@@ 4.1 Mean Absolute Error (MAE)
Mean Absolute Error (MAE) is a widely used metric for assessing the accuracy of predictions in various applications, including video prediction [51]. MAE quantifies the average magnitude of errors between predicted and actual pixel positions of video frames, without considering their direction. MAE is calculated as:
@@ 4.2 JEPA Embedding Distance (JEDi)
The JEPA (Joint Embedding Predictive Architecture) Embedding Distance (JEDi) [30] offers a powerful approach for evaluating video generation quality by measuring the distributional similarity between generated and real video sequences. Addressing limitations of metrics like Fréchet Video Distance (FVD) [56], which assumes Gaussianity of the underlying video data, JEDi does not require such distributional assumptions and leverages Maximum Mean Discrepancy (MMD) with a polynomial kernel to compute the distance between JEPA-derived video feature sets. MMD is an integral probability metric which quantifies the difference between two probability distributions by finding a function within a chosen class that maximizes the difference in expectations over the distributions [17]. In addition to its applicability over a broad range of distributions, JEDi also aligns more closely with human evaluations compared to FVD and offers a scalable alternative to manually collecting human opinion scores, making it well-suited for large-scale evaluation of generative video models. The setup for JEDi estimation is described briefly as below.
@@ 4.3 Average Pair-wise Displacement (APD)
Average Pair-wise Displacement (APD) [5] is designed to evaluate the diversity of predicted trajectories in multi-modal forecasting scenarios. In this work we use APD to measure diversity across videos generated using our real-time motion transfer pipeline. Unlike traditional metrics that compare each prediction to the ground truth, APD measures diversity by computing the average pairwise displacement across all predicted trajectories for a given video. In the context of diverse video prediction, it quantifies how much the generated video samples differ from one another over a time interval. APD is defined as:
@@
where $M$ is the number of generated samples for a given video, P is the number of pixel values in a video frame, and ${}^{(i)}\hat{\mathbf{v}}_{t_{1}:t_{2}}^{a}$ and ${}^{(j)}\hat{\mathbf{v}}_{t_{1}:t_{2}}^{a}$ denote the $i$-th and $j$-th predicted trajectory for a video $a$ over the time interval from $t_{1}$ to $t_{2}$, respectively. The numerator sums the Manhattan distances between every pair of predicted trajectories, while the denominator normalizes this total by the number of prediction pairs, the length of the prediction horizon and the number of pixel values in a frame. A higher APD value indicates that the predictions are more widely dispersed, reflecting a model’s ability to generate diverse future trajectories rather than merely converging on the most likely outcome. An APD value which is close to 00 indicates that there is almost no diversity among the generated samples.
@@ 4.4 Structural Similarity (SSIM)
The Structural Similarity Index (SSIM) [64] is a quality metric which is used for evaluating the perceptual quality of generated images. It quantifies the similarity between a generated frame and the corresponding ground truth frame by jointly comparing luminance, contrast, and structural information. It has been proven that for estimating video reconstruction quality from the perspective of the human visual system, SSIM is a more effective metric than the commonly used Peak Signal-to-Noise Ratio (PSNR) [23]. The SSIM metrics lies between 00 and $1$ where a value of 00 indicates no similarity between generated video frame and ground truth video frame, and $1$ indicates 100% similarity between the two.
@@
where g is the generated video frame and r is the real video frame, $\mu_{g}$ and $\mu_{r}$ represent the mean pixel intensities of frames $g$ and $r$, $\sigma_{g}^{2}$ and $\sigma_{r}^{2}$ denote their variances, and $\sigma_{gr}$ is the covariance between the two frames. The constants $c_{1}$ and $c_{2}$ are included to stabilize the division in cases of denominator values close to 00.
5 Results and Discussion
@@ 5.1 Datasets
Mgif dataset: We employ 284 videos for training and 34 for testing from the MGIF dataset with 256 x 256 resolution [50], which records periodic animal movements. These videos exhibit a range of frame counts, from 168 to 10,500 frames and allows for a thorough assessment of the model’s capability to recognize distinct movement patterns among various animal species.
Bair dataset: For the BAIR dataset [12], we work with 5001 training videos and 256 testing videos with 256 x 256 resolution, each consisting of 30 frames. These depict robotic arms interacting with different objects, offering a rigorous test of the model’s effectiveness in understanding dynamic environments with complex backgrounds and irregular motions.
VoxCeleb dataset: We utilize 3884 training videos and 44 testing videos from the VoxCeleb dataset with 256 x 256 resolution [36], which includes interview clips featuring different celebrities. The video lengths vary with frame counts ranging from 72 to 1228 frames. This dataset is particularly valuable for assessing the model’s proficiency in picking up on subtle facial expressions which is a critical requirement for applications like video conferencing where conveying nuanced emotions is important.
@@ 5.2 Single Video Sample Prediction
We have applied the GRU, VAE, VRNN, NF and GRU-NF models to the three datasets in both reconstruction and transfer modes for single sample prediction. The prediction is performed using two different prediction horizons in order to observe how the models perform over varying horizon lengths. For each dataset the prediction horizons are generated using nonoverlapping sliding windows over the dataset. For both the Mgif and VoxCeleb datasets, we use horizon lengths of 6 and 12 where the first 6 frames are used as input and next 6 frames as output, and first 12 frames are used as input and next 12 frames as output respectively. For the Bair dataset which has a fixed set of 30 frames for each video, we set the prediction horizon to 5 and 15 and in this case we use the first 5 frames as input and next 5 frames as output, and the first 15 frames as input and next 15 frames as output respectively.
@@
For assessing single sample prediction in reconstruction and transfer modes, we have used the MAE and JEDi metrics. As there is no original video for transfer mode, ground truth videos are generated using the FOMM piepline with the original keypoints of the driving video and source image. The results for both reconstruction and transfer mode for Mgif, Bair and VoxCeleb dataset are demonstrated in Tables 1, 2, 3, 4, 5 and 6. The first column of the tables represents the number of input frames and the number of output frames used for prediction. From these results it can be seen that the VRNN significantly outperforms all other models. In terms of MAE, VRNN consistently achieves the lowest values across all models for each dataset, and for each mode and each prediction horizon. For JEDi, VRNN is the best-performing model in 8 out of 12 cases, while GRU and GRU-NF achieve the best performance in the remaining 4 cases. This trend holds across all datasets where VRNN demonstrates robust performance.
@@
Qualitative results using frames of generated videos are also shown for VoxCeleb data for both reconstruction and transfer modes and 12-12 prediction horizon in Figures 8 and 9 respectively for single sample prediction. As reference the second row of the figure shows the video frames that are generated using the FOMM pipeline without keypoint prediction. For reconstruction mode, driving video sequence shown in the first row is the ground truth video and for transfer mode, FOMM generated video sequence is the ground truth. The third, fourth, fifth, sixth and seventh rows show the video frames generated using VRNN, VAE, GRU, NF and GRU-NF respectively for keypoint prediction and FOMM pipeline. It is evident from the figures that the output using VRNN is almost similar to the ground truth, unlike the generated outputs from the other models.
@@
Table 1: MAE and JEDi results of Mgif dataset on reconstruction mode
Prediction range MAE JEDi
VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
6 - 6 0.0338 0.0392 0.0373 0.0379 0.0376 6.4237 6.3117 6.4354 6.2996 6.2336
12 -12 0.0338 0.0388 0.0379 0.0379 0.0377 6.5757 6.5748 6.4593 6.085 6.0207
@@
Table 2: MAE and JEDi results of Mgif dataset on transfer mode
Prediction range MAE JEDi
VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
6 - 6 0.016 0.0215 0.0195 0.0201 0.0197 0.2738 0.9997 0.4312 1.0332 0.9396
12 - 12 0.0156 0.0215 0.0199 0.0202 0.0201 0.2821 0.464 0.423 1.2005 0.8449
@@
Table 3: MAE and JEDi results of Bair dataset on reconstruction mode
Prediction range MAE JEDi
VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
5 - 5 0.0664 0.0749 0.0722 0.0766 0.0763 1.7068 2.3094 1.6094 2.5065 2.7014
15 - 15 0.0715 0.0759 0.0773 0.0793 0.0791 2.4425 3.7483 3.9859 3.4577 3.665
@@
Table 4: MAE and JEDi results of Bair dataset on transfer mode
Prediction range MAE JEDi
VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
5 - 5 0.0139 0.0267 0.0216 0.0272 0.0267 0.5898 1.5357 0.5262 1.5638 1.9548
15 - 15 0.0203 0.0263 0.0277 0.03 0.0298 0.9373 3.8864 3.5842 3.1205 2.8107
@@
Table 5: MAE and JEDi results of VoxCeleb dataset on reconstruction mode
Prediction range MAE JEDi
VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
6 - 6 0.0478 0.0602 0.0514 0.0729 0.0715 0.6579 0.8189 0.6936 2.3025 2.2377
12 - 12 0.0472 0.0651 0.0566 0.0724 0.0715 0.666 1.2794 0.7302 1.4351 1.6561
@@
Table 6: MAE and JEDi results of VoxCeleb dataset on transfer mode
Prediction range MAE JEDi
VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
6 - 6 0.0091 0.0244 0.0148 0.0364 0.0348 0.5166 0.5583 0.5808 2.4871 2.0271
12 - 12 0.0077 0.0287 0.0205 0.035 0.0348 0.5448 0.7610 0.6743 1.1633 1.2116
@@ 5.3 Diverse Video Sample Prediction
We have conducted our experiments using VRNN and GRU-NF for diverse sample prediction using the longer horizon, i.e. the 12-12 prediction range and generated 100 samples for each test video. This is done in both reconstruction and transfer modes for the VoxCeleb and Bair datasets. For this purpose we first select $C=20$ video samples from the $D=100$ generated samples that have the lowest MAE values compared to ground truth to demonstrate the diversity-fidelity tradeoff. Here the values C and D have been selected to compare the diverse video generation capabilities of the two models and in general they will vary based on the given application. To assess the performance of diverse sample generation, we have used APD to measure the pairwise distance among these 20 generated samples. The model that produces more diverse samples will have a higher APD value, indicating that the samples differ from one another significantly. Following this, the ratio of APD to MAE of each model is calculated to evaluate diversity vs. fidelity. A higher APD-to-MAE ratio for a given model indicates that it can produce more diverse samples with an acceptable difference wrt the ground truth. To assess perceptual similarity of the generated videos wrt the ground truth we have also measured the SSIM values of the generated video frames.
Table 7: Diversity vs. fidelity results of VoxCeleb dataset on reconstruction mode
APD MAE SSIM APD to MAE ratio
Mean Median
VRNN 0.009 0.0533 0.7428 0.3014 0.2298
GRU-NF 0.0407 0.0817 0.6235 1.1871 1.2153
@@
Here, $APD_{\min}$ and $APD_{\max}$, and $MAE_{\min}$ and $MAE_{\max}$ represent the minimum and maximum values of the APD and MAE values respectively. These are determined using the complete set of APD and MAE values across all test videos and both models to ensure a consistent normalization scale facilitating direct comparisons. After standardizing the APD and MAE values, the ratio of standardized APD to standardized MAE is calculated for each test video. The mean and median of these ratio values across all test videos are then reported in Tables 7, 8, 9 and 10. Since APD and MAE represent different performance metrics, our standardization procedure ensures that their values remain within the same range, thereby allowing for a meaningful comparison of their ratios.
@@
In Table 7, diverse sample prediction performance comparisons of VRNN and GRU-NF models are given for VoxCeleb dataset in reconstruction mode. It shows that GRU-NF produces more diverse samples compared to VRNN since the APD value of GRU-NF is more than 4.52 times higher than that of VRNN. Although the MAE value of VRNN is lower than that of GRU-NF, the APD to MAE ratio of GRU-NF indicates that the generated diverse samples in this case do not significantly deviate from the ground truth. The mean value of the APD to MAE ratios for GRU-NF is 3.94 times higher than that of VRNN, indicating that although the samples generated from VRNN’s latent space are close to the ground truth, they lack significant diversity. The median value of the ratios also reflects this trend, with a higher value for GRU-NF compared to VRNN. This suggests that, even across different test videos, GRU-NF consistently produces a diverse set of predictions with an acceptable deviation from ground truth. A higher median value observed in GRU-NF further reinforces that the higher APD to MAE ratio relative to VRNN is not due to a few outliers but a general characteristic of the model across the dataset. A similar trend is also seen in the results as given in Tables 8, 9 and 10. In each case, the mean and median values of APD to MAE ratios are higher for GRU-NF than VRNN.
@@
To provide a comparison of the standardized APD to standardized MAE ratio distributions for both models, we use kernel density plots shown in Figures 13 and 13 for VoxCeleb dataset in reconstruction mode and transfer mode respectively. Similarly, the comparison for Bair dataset in both modes for both models are shown in Figures 13 and 13. To summarize, across both datasets and both modes, the pattern remains consistent: the density peak of GRU-NF appears to the right of VRNN’s peak, meaning it prioritizes diversity but not at the cost of excessive error. Conversely, the density distributions of VRNN reflect slightly lower error along with lower diversity. Therefore, GRU-NF offers a better and reasonable tradeoff between fidelity and diversity.
@@
To assess the perceptual quality of the generated samples, we also perform a qualitative evaluation based on structural similarity. Specifically, for each test video, we select the 20 generated samples with the lowest MAE values and compute the SSIM between each of these samples and the corresponding ground truth. The mean of the SSIM values are then taken to obtain a single SSIM score per video. Finally, the mean SSIM value across all test videos is reported in the Tables 7, 8, 9 and 10. The results in Table 7 indicate that the SSIM value is greater for VRNN than GRU-NF since the generated samples using VRNN have little deviation from ground truth compared to GRU-NF. However, the samples generated using GRU-NF have achieved 62.35% SSIM which means that the video frame quality can be characterized as good [44]. Similarly in Tables 8, 9 and 10, the SSIM values indicate that though VRNN has greater SSIM values than GRU-NF, SSIM values for GRU-NF fall in good or excellent quality range [44]. Thus, the results of the APD to MAE ratio and SSIM value across both modes of both datasets demonstrate that the generated samples using GRU-NF are more diverse compared to VRNN, while having meaningful representations.
@@
Qualitative results for diverse sample prediction are shown in Figures 14 and 15 for VoxCeleb dataset. In Figure 14, the upper panel and the lower panel show the produced diverse samples in reconstruction mode using GRU-NF and VRNN for keypoint prediction respectively, and in Figure 15, the upper panel and the lower panel show the produced diverse samples in transfer mode using GRU-NF and VRNN for keypoint prediction respectively. In each figure, the first row serves as the ground truth video sequence. The second, third and fourth rows show the examples of generated diverse video frames using the prediction model. It is evident from the figures that the facial expressions in the generated video samples using VRNN for keypoint prediction are nearly identical. In contrast, the generated samples using GRU-NF for keypoint prediction exhibit distinct facial expressions, demonstrating greater variation while remaining meaningful.

### v2

@@ 4.1 Mean Absolute Error (MAE)
Mean Absolute Error (MAE) is a widely used metric for assessing the accuracy of predictions in various applications, including video prediction [siarohin2019first]. MAE quantifies the average magnitude of errors between predicted and actual pixel positions of video frames, without considering their direction. MAE is calculated as:
@@ 4.2 JEPA Embedding Distance (JEDi)
The JEPA (Joint Embedding Predictive Architecture) Embedding Distance (JEDi) [luo2024beyond] offers a powerful approach for evaluating video generation quality by measuring the distributional similarity between generated and real video sequences. Addressing limitations of metrics like Fréchet Video Distance (FVD) [unterthiner2018fvd], which assumes Gaussianity of the underlying video data, JEDi does not require such distributional assumptions and leverages Maximum Mean Discrepancy (MMD) with a polynomial kernel to compute the distance between JEPA-derived video feature sets. MMD is an integral probability metric which quantifies the difference between two probability distributions by finding a function within a chosen class that maximizes the difference in expectations over the distributions [gretton2012kernel]. In addition to its applicability over a broad range of distributions, JEDi also aligns more closely with human evaluations compared to FVD and offers a scalable alternative to manually collecting human opinion scores, making it well-suited for large-scale evaluation of generative video models. The setup for JEDi estimation is described briefly as below.
@@ 4.3 Average Pair-wise Displacement (APD)
Average Pair-wise Displacement (APD) [chen2024mixed] is designed to evaluate the diversity of predicted trajectories in multi-modal forecasting scenarios. In this work we use APD to measure diversity across videos generated using our real-time motion transfer pipeline. Unlike traditional metrics that compare each prediction to the ground truth, APD measures diversity by computing the average pairwise displacement across all predicted trajectories for a given video. In the context of diverse video prediction, it quantifies how much the generated video samples differ from one another over a time interval. APD is defined as:
@@
where $M$ is the number of generated samples for a given video, P is the number of pixel values in a video frame, and $^{(i)}\hat{\mathbf{v}}_{t_{1}:t_{2}}^{a}$ and $^{(j)}\hat{\mathbf{v}}_{t_{1}:t_{2}}^{a}$ denote the $i$-th and $j$-th predicted trajectory for a video $a$ over the time interval from $t_{1}$ to $t_{2}$, respectively. The numerator sums the Manhattan distances between every pair of predicted trajectories, while the denominator normalizes this total by the number of prediction pairs, the length of the prediction horizon and the number of pixel values in a frame. A higher APD value indicates that the predictions are more widely dispersed, reflecting a model’s ability to generate diverse future trajectories rather than merely converging on the most likely outcome. An APD value which is close to 0 indicates that there is almost no diversity among the generated samples.
@@ 4.4 Structural Similarity (SSIM)
The Structural Similarity Index (SSIM) [Wang2004SSIM] is a quality metric which is used for evaluating the perceptual quality of generated images. It quantifies the similarity between a generated frame and the corresponding ground truth frame by jointly comparing luminance, contrast, and structural information. It has been proven that for estimating video reconstruction quality from the perspective of the human visual system, SSIM is a more effective metric than the commonly used Peak Signal-to-Noise Ratio (PSNR) [Kotevski2009]. The SSIM metrics lies between 0 and 11 where a value of 0 indicates no similarity between generated video frame and ground truth video frame, and 11 indicates 100% similarity between the two.
@@
where g is the generated video frame and r is the real video frame, $\mu_{g}$ and $\mu_{r}$ represent the mean pixel intensities of frames $g$ and $r$, $\sigma_{g}^{2}$ and $\sigma_{r}^{2}$ denote their variances, and $\sigma_{gr}$ is the covariance between the two frames. The constants $c_{1}$ and $c_{2}$ are included to stabilize the division in cases of denominator values close to 0.
5 Results
@@ 5.1 Datasets
MGIF dataset: We employ 284 videos for training and 34 for testing from the MGIF dataset with 256 x 256 resolution [siarohin2019animating], which records periodic animal movements. These videos exhibit a range of frame counts, from 168 to 10,500 frames and allow for a thorough assessment of the model’s capability to recognize distinct movement patterns among various animal species.
BAIR dataset: For the BAIR dataset [ebert2017self], we work with 5001 training videos and 256 testing videos with 256 x 256 resolution, each consisting of 30 frames. These depict robotic arms interacting with different objects, offering a rigorous test of the model’s effectiveness in understanding dynamic environments with complex backgrounds and irregular motions.
VoxCeleb dataset: We utilize 3884 training videos and 44 testing videos from the VoxCeleb dataset with 256 x 256 resolution [nagrani2017voxceleb], which includes interview clips featuring different celebrities. The video lengths vary with frame counts ranging from 72 to 1228 frames. This dataset is particularly valuable for assessing the model’s proficiency in picking up on subtle facial expressions which is a critical requirement for applications like video conferencing where conveying nuanced emotions is important.
The MGIF and BAIR datasets serve as our non-human benchmarks. MGIF captures motion of deformable objects and animated characters, while BAIR consists of robotic arm manipulation sequences. Including these datasets allows us to evaluate the proposed models under non-human and highly deformable motion regimes, complementing the human-centric VoxCeleb dataset.
@@ 5.2 Single Video Sample Prediction
We have applied the GRU, VAE, VRNN, NF and GRU-NF models to the three datasets in both reconstruction and transfer modes for single sample prediction. The prediction is performed using three different prediction horizons in order to observe how the models perform over varying horizon lengths. For each dataset the prediction horizons are generated using nonoverlapping sliding windows over the dataset. For both the MGIF and VoxCeleb datasets, we evaluate three prediction settings: 6–6, 6–24, and 12–12. In 6–6 and 6-24, the first 6 frames are used as input and the subsequent 6 frames or 24 frames are predicted respectively. In 12–12, the first 12 frames are used as input and the subsequent 12 frames are predicted. For the BAIR dataset, we evaluate 5–5, 5–20, and 15–15. In 5–5 and 5-20, the first 5 frames are used as input and the next 5 frames or 20 frames are predicted respectively. In 15–15, the first 15 frames are used as input and the next 15 frames are predicted.
@@
For assessing single sample prediction in reconstruction and transfer modes, we have used the MAE and JEDi metrics. As there is no original video for transfer mode, ground truth videos are generated using the FOMM pipeline with the original keypoints of the driving video and source image. The results for both reconstruction and transfer mode for MGIF, BAIR and VoxCeleb dataset are demonstrated in Tables 2, 2, 4, 4, 6 and 6. The first column of the tables represents the number of input frames and the number of output frames used for prediction.
Overall, the results show a systematic dependence on dataset characteristics and prediction horizons, rather than a single model dominating all settings. In terms of pixel-level fidelity, VRNN achieves the best performance in the majority of cases (13 out of 18 settings), indicating strong reconstruction accuracy across datasets. Conversely, the JEDi metric which reflects spatiotemporal coherence tends to favor GRU. This implies that deterministic recurrent models can better match global temporal statistics under low entropy motion conditions.
This behavior is consistent with the conditional multi-modality of the datasets and the forecast horizons used in our study. In case of BAIR, the VRNN attains the best JEDi at longer horizons e.g., 5–20 and 15–15 in both reconstruction and transfer modes, indicating that latent variable dynamics better capture multi-modal future evolution as the prediction horizon increases. In contrast GRU is favored on MGIF/VoxCeleb where the motion is more repeatable or primarily involves subtle facial changes, and deterministic models are sufficient to match global temporal feature statistics.
@@
Qualitative results using frames of generated videos are also shown for VoxCeleb data for both reconstruction and transfer modes and 12-12 prediction horizon in Figures 8 and 9 respectively for single sample prediction. As reference the second row of the figure shows the video frames that are generated using the FOMM pipeline without keypoint prediction. For reconstruction mode, driving video sequence shown in the first row is the ground truth video and for transfer mode, FOMM generated video sequence is the ground truth. The third, fourth, fifth, sixth and seventh rows show the video frames generated using VRNN, VAE, GRU, NF and GRU-NF respectively for keypoint prediction and FOMM pipeline.
@@
Table 1: MAE and JEDi results of MGIF dataset on reconstruction mode. For each criteria i.e. MAE and JEDi the best model is highlighted in bold, lower the better.
Prediction MAE JEDi
range VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
6–6 0.0316 0.0392 0.0373 0.0379 0.0376 6.3555 6.3117 6.4354 6.2996 6.2336
6–24 0.0471 0.0501 0.0451 0.0467 0.0454 6.5361 7.4836 6.1183 6.4495 6.7648
12–12 0.0375 0.0388 0.0379 0.0379 0.0377 5.8972 6.5748 6.4593 6.0850 6.0207
@@
Table 2: MAE and JEDi results of MGIF dataset on transfer mode. For each criteria i.e. MAE and JEDi the best model is highlighted in bold, lower the better.
Prediction MAE JEDi
range VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
6–6 0.0192 0.0215 0.0195 0.0201 0.0197 0.6635 0.9997 0.4312 1.0332 0.9396
6–24 0.0343 0.0385 0.0319 0.0339 0.0321 0.8436 5.2887 0.695 1.0901 1.2615
12–12 0.0199 0.0215 0.0199 0.0202 0.0201 0.983 0.464 0.423 1.2005 0.8449
@@
Table 3: MAE and JEDi results of BAIR dataset on reconstruction mode. For each criteria i.e. MAE and JEDi the best model is highlighted in bold, lower the better.
Prediction MAE JEDi
range VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
5–5 0.0713 0.0749 0.0722 0.0766 0.0763 1.655 2.3094 1.6094 2.5065 2.7014
5–20 0.0891 0.0911 0.09 0.0911 0.091 6.2676 6.8537 7.8202 7.1344 8.3909
15–15 0.0765 0.0759 0.0773 0.0793 0.0791 3.2155 3.7483 3.9859 3.4577 3.665
@@
Table 4: MAE and JEDi results of BAIR dataset on transfer mode. For each criteria i.e. MAE and JEDi the best model is highlighted in bold, lower the better.
Prediction MAE JEDi
range VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
5–5 0.0199 0.0267 0.0216 0.0272 0.0267 0.9274 1.5357 0.5262 1.5638 1.9548
5–20 0.0449 0.0468 0.0456 0.0468 0.0467 5.7885 6.8148 7.9668 7.1783 7.734
15–15 0.0258 0.0263 0.0277 0.03 0.0298 1.8371 3.8864 3.5842 3.1205 2.8107
@@
Table 5: MAE and JEDi results of VoxCeleb dataset on reconstruction mode. For each criteria i.e. MAE and JEDi the best model is highlighted in bold, lower the better.
Prediction MAE JEDi
range VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
6–6 0.0511 0.0602 0.0514 0.0729 0.0715 0.7079 0.8189 0.6936 2.3025 2.2377
6–24 0.0707 0.0771 0.0732 0.0875 0.0878 1.105 0.9118 0.9462 3.0365 3.3046
12–12 0.0564 0.0651 0.0566 0.0724 0.0715 0.7664 1.2794 0.7302 1.4351 1.6561
@@
Table 6: MAE and JEDi results of VoxCeleb dataset on transfer mode. For each criteria i.e. MAE and JEDi the best model is highlighted in bold, lower the better.
Prediction MAE JEDi
range VRNN VAE GRU NF GRU-NF VRNN VAE GRU NF GRU-NF
6–6 0.0137 0.0244 0.0148 0.0364 0.0348 0.5918 0.5583 0.5808 2.4871 2.0271
6–24 0.0442 0.0445 0.0411 0.0539 0.0553 1.2844 0.8407 0.9875 3.4093 2.8105
12–12 0.0203 0.0287 0.0205 0.035 0.0348 0.8267 0.7610 0.6743 1.1633 1.2116
@@ 5.3 Diverse Video Sample Prediction
We have conducted our experiments using VRNN and GRU-NF for diverse sample prediction at longer prediction horizons, i.e. 6–24 and 12–12 for VoxCeleb, and 5–20 and 15–15 for BAIR. For each test video, we generated 100 samples in both reconstruction and transfer modes. We then select $C=20$ video samples from the $D=100$ generated samples that have the lowest MAE values compared to ground truth to demonstrate the diversity-fidelity trade-off. Here the values C and D have been selected to compare the diverse video generation capabilities of the two models and in general they will vary based on the given application. To assess the performance of diverse sample generation, we have used APD to measure the pairwise distance among these 20 generated samples. The model that produces more diverse samples will have a higher APD value, indicating that the samples differ from one another significantly. Following this, the ratio of APD to MAE of each model is calculated to evaluate diversity vs. fidelity. A higher APD-to-MAE ratio for a given model indicates that it can produce more diverse samples with an acceptable difference w.r.t. the ground truth. To assess perceptual similarity of the generated videos w.r.t. the ground truth we have also measured the SSIM values of the generated video frames.
Table 7: Diversity vs. fidelity results of VoxCeleb dataset on reconstruction mode. For APD to MAE ratio the best model is highlighted in bold, higher the better.
Prediction Model APD MAE SSIM APD to MAE ratio
range Mean Median
12 - 12 VRNN 0.0074 0.062 0.6914 0.1142 0.0876
GRU-NF 0.0407 0.0817 0.6235 1.3452 1.3254
6 - 24 VRNN 0.0079 0.0814 0.6026 0.0819 0.0749
GRU-NF 0.0608 0.0999 0.5441 1.6321 1.3126
Table 8: Diversity vs. fidelity results of VoxCeleb dataset on transfer mode. For APD to MAE ratio the best model is highlighted in bold, higher the better.
Prediction Model APD MAE SSIM APD to MAE ratio
range Mean Median
12 - 12 VRNN 0.007 0.0334 0.8497 1.268 1.0229
GRU-NF 0.0336 0.0464 0.7861 5.4636 5.3986
6 - 24 VRNN 0.0074 0.0548 0.7306 0.3199 0.3217
GRU-NF 0.05 0.0701 0.6655 3.9243 3.6426
Table 9: Diversity vs. fidelity results of BAIR dataset on reconstruction mode. For APD to MAE ratio the best model is highlighted in bold, higher the better.
Prediction Model APD MAE SSIM APD to MAE ratio
range Mean Median
15 - 15 VRNN 0.0113 0.0834 0.7413 0.4445 0.3318
GRU-NF 0.0269 0.0853 0.7377 1.7144 1.369
5 - 20 VRNN 0.0103 0.096 0.714 0.2831 0.2355
GRU-NF 0.0253 0.0971 0.7106 1.095 0.9684
Table 10: Diversity vs. fidelity results of BAIR dataset on transfer mode. For APD to MAE ratio the best model is highlighted in bold, higher the better.
Prediction Model APD MAE SSIM APD to MAE ratio
range Mean Median
15 - 15 VRNN 0.011 0.0364 0.9065 0.519 0.3444
GRU-NF 0.0268 0.0393 0.8996 1.4679 1.4192
5 - 20 VRNN 0.0099 0.053 0.8609 0.339 0.2606
GRU-NF 0.0247 0.0541 0.857 1.2689 1.1047
@@
Here, $APD_{\min}$ and $APD_{\max}$, and $MAE_{\min}$ and $MAE_{\max}$ represent the minimum and maximum values of the APD and MAE values respectively. These are determined using the complete set of APD and MAE values across all test videos and both models to ensure a consistent normalization scale facilitating direct comparisons. After standardizing the APD and MAE values, the ratio of standardized APD to standardized MAE is calculated for each test video. The mean and median of these ratio values across all test videos are then reported in Tables 7, 8, 10 and 10. Since APD and MAE represent different performance metrics, our standardization procedure ensures that their values remain within the same range, thereby allowing for a meaningful comparison of their ratios.
@@
In Table 7, diverse sample prediction performance comparisons of VRNN and GRU-NF models are given for VoxCeleb dataset in reconstruction mode for 6-24 and 12-12 prediction range. It shows that in 12-12 prediction horizon GRU-NF produces more diverse samples compared to VRNN since the APD value of GRU-NF is more than 5.5 times higher than that of VRNN. Although the MAE value of VRNN is lower than that of GRU-NF, the APD to MAE ratio of GRU-NF indicates that the generated diverse samples in this case do not significantly deviate from the ground truth. The mean value of the APD to MAE ratios for GRU-NF is 11.78 times higher than that of VRNN, indicating that although the samples generated from VRNN’s latent space are close to the ground truth, they lack significant diversity. The median value of the ratios also reflects this trend, with a higher value for GRU-NF compared to VRNN. This suggests that, even across different test videos, GRU-NF consistently produces a diverse set of predictions with an acceptable deviation from ground truth. A higher median value observed in GRU-NF further reinforces that the higher APD to MAE ratio relative to VRNN is not due to a few outliers but a general characteristic of the model across the dataset. This trend can also be observed for the 6-24 prediction horizon. A similar pattern is seen in the results as given in Tables 8, 10 and 10. In each case, the mean and median values of APD to MAE ratios are higher for GRU-NF than VRNN.
@@
To provide a comparison of the standardized APD to standardized MAE ratio distributions for both models, we use kernel density plots shown in Figures 10 and 11 for VoxCeleb dataset in reconstruction mode and transfer mode respectively for 12-12 prediction horizon. Similarly, the comparison for BAIR dataset in both modes for both models for 15-15 prediction horizon are shown in Figures 12 and 13 for 15-15 prediction horizon. To summarize, across both datasets and both modes, the pattern remains consistent: the density peak of GRU-NF appears to the right of VRNN’s peak, meaning it prioritizes diversity but not at the cost of excessive error. Conversely, the density distributions of VRNN reflect slightly lower error along with lower diversity. Therefore, GRU-NF offers a better and reasonable trade-off between fidelity and diversity.
@@
To assess the perceptual quality of the generated samples, we also perform a qualitative evaluation based on structural similarity. Specifically, for each test video, we select the 20 generated samples with the lowest MAE values and compute the SSIM between each of these samples and the corresponding ground truth. The mean of the SSIM values are then taken to obtain a single SSIM score per video. Finally, the mean SSIM value across all test videos is reported in the Tables 7, 8, 10 and 10. The results indicate that the SSIM value is greater for VRNN than GRU-NF since the generated samples using VRNN have little deviation from ground truth compared to GRU-NF. However though VRNN has greater SSIM values than GRU-NF, SSIM values for GRU-NF fall in fair, good or excellent quality range [Rehman2015]. Thus, the results of the APD to MAE ratio and SSIM value across both modes of both datasets demonstrate that the generated samples using GRU-NF are more diverse compared to VRNN, while having meaningful representations.
@@
Qualitative results for diverse sample prediction are shown in Figures 14 and 15 for VoxCeleb dataset for 12-12 prediction horizon. The Figures 14(a) and 14(b) show the produced diverse samples in reconstruction mode using GRU-NF and VRNN for keypoint prediction respectively, and Figures 15(a) and 15(b) show the produced diverse samples in transfer mode using GRU-NF and VRNN for keypoint prediction respectively. In each figure, the first row serves as the ground truth video sequence. The second, third and fourth rows show the examples of generated diverse video frames using the prediction model. It is evident from the figures that the facial expressions in the generated video samples using VRNN for keypoint prediction are nearly identical. In contrast, the generated samples using GRU-NF for keypoint prediction exhibit distinct facial expressions, demonstrating greater variation while remaining meaningful.